A man walks 240m in 30s & then walk back 140m toward the starting point in the next 35s. What is the average speed and velocity?
1
Expert's answer
2011-08-26T06:34:44-0400
Recall that the speed is the absolute value of velocity.The average speed is
(240 +140)/(35+30)= 380 / 65 = 5.85 m/s

Call the direction of the first 30s move "positive".
Since during next 35s the man moves in "negative" direction, the average velocity is
(240 - 140)/(35+30)= 100 / 65 = 1.54 m/s in positive direction.